# Interpretable Plant Leaf Disease Detection Using Attention-Enhanced CNN

### v1

@@ Abstract
Plant diseases pose a significant threat to global food security, necessitating accurate and interpretable disease detection methods. This study introduces an interpretable attention-guided Convolutional Neural Network (CNN), CBAM-VGG16, for plant leaf disease detection. By integrating Convolution Block Attention Module (CBAM) at each convolutional stage, the model enhances feature extraction and disease localization. Trained on five diverse plant disease datasets, our approach outperforms recent techniques, achieving high accuracy (up to 98.87%) and demonstrating robust generalization. Here, we show the effectiveness of our method through comprehensive evaluation and interpretability analysis using CBAM attention maps, Grad-CAM, Grad-CAM++, and Layer-wise Relevance Propagation (LRP). This study advances the application of explainable AI in agricultural diagnostics, offering a transparent and reliable system for smart farming. The code of our proposed work is available at https://github.com/BS0111/PlantAttentionCBAM
@@
Agriculture, Plant leaf disease, Artificial intelligence, Deep learning, Attention, Interpretability
@@ 1 Introduction
The agriculture industry is essential in maintaining food security worldwide, but different crops suffer from a variety of diseases due to varying weather conditions, posing a major threat to crop yield and quality. These diseases are often caused by factors such as extreme temperatures, microbial infections, and changes in humidity or soil conditions. Farmers have always relied on human examination to identify diseases, which is labour-intensive, prone to mistakes, and ineffective on large farms. As a result, among the most crucial areas of research in smart agriculture is the automation of plant disease identification and classification utilizing various Artificial Intelligence (AI) approaches [Xiang2021]. Machine learning methods in agriculture has drawn a lot of interest with applications ranging from yield prediction and crop recommendation to disease detection. In recent years, Convolutional Neural Networks (CNNs) and transformers have shown impressive performance in a variety of computer vision applications such as segmentation, object recognition, and image denoising [Lin2025SCNet]. CNN-based models in agriculture have performed well on challenges such as classifying plant diseases. However, despite their effectiveness, CNNs often operate as black boxes, offering limited transparency into how predictions are made, hampering trust and broader adoption.
@@
In order to tackle this, Explainable AI (XAI) methods have been developed enabling visual interpretation of model decisions through attention maps [LinEAPT2023], GradCAM [Selvaraju2017], and LRP [Bach2015]. While GradCAM and GradCAM++ [Chattopadhay2018] generate class-discriminative localization maps using gradients from the final convolutional layers, they may lack resolution and be susceptible to noisy activations. LRP, in contrast, provides pixel-level attributions by backpropagating the model output through a set of layer-specific relevance propagation rules. This offers a more granular explanation of predictions, which is crucial in medical and agricultural diagnostics.
@@
In this work, an explainable deep learning approach is proposed for the detection of plant leaf disease. Our architecture is based on the VGG16 [Simonyan2014] backbone enhanced with the CBAM [woo2018cbamconvolutionalblockattention, Labiadh2025] which introduces attention layers for inherent interpretability of the model with an emphasis on the most relevant features at both the channel and spatial levels. Following each of the five convolutional layers, CBAM modules are added to improve classification accuracy and localization of relevant features by capturing both spatial and channel-wise attention. Five distinct datasets are used to train the model, namely Apple, Plant Village, Embrapa, Maize and Rice to ensure the generlaizability and applicability of our proposed method across diverse set of crops. Apart from the inherent interpretability of the proposed method’s decision-making process through CBAM layers, we also demonstrate interpretability using advanced explainability methods like LRP, Grad-CAM, and Grad-CAM++. We have also employed high-dimensional feature visualization techniques such as t-distributed Stochastic Neighbor Embedding (t-SNE) and Uniform Manifold Approximation and Projection (UMAP) to visualize feature in lower dimension for visualization of extracted features. Overall, this study advances the application of XAI for agricultural use by presenting an interpretable and performance-robust framework for plant disease classification. The following is a summary of the main contributions of the proposed work:
@@
  • The proposed method demonstrates better performance compared to other recent cutting-edge techniques on all five datasets considered in this work ensuring generalization and adaptive capability of proposed method for any crop leaf disease identification.
In rest of the paper, Section 2 reviews existing literature, highlighting key limitations and motivations. Section 3 details the architecture of the proposed CBAM-VGG16 model, including the integration of attention mechanisms, and interpretability approaches. Section 4 describes the complete results including the experimental setup, dataset description, comparative results, and the explainability analysis. Finally, Section 5 brings the study to a conclusion and suggests possible directions for further research in this domain.
@@ 2 Related Works
Deep learning combined with image processing has been frequently utilized for the automatic detection of diseases in plant leaves. Numerous Research has explored various models, datasets, and image processing pipelines for this task. This section offers a thorough analysis of the methods employed for multiclass leaf disease classification, emphasizing traditional techniques, deep learning methods, and recent advances in vision transformers [Zhang2025, ZHANGTransformer2023].
[Mohanty2016] put together a a large dataset known as plant village that included 38 distinct pictures of crop leaf diseases. They assessed the dataset using several popular deep learning frameworks like AlexNet and GoogleNet, employing a transfer learning approach. [Barbedo2018] applied conventional Convolutional Neural Network (CNN) architectures to analyze how variations in samples and dataset size influence The efficiency of identifying plant diseases via transfer learning. For this purpose, a dataset comprising 1,383 background-free images across 56 distinct disease types was compiled. [Singh2019] developed a Multilayer Convolutional Neural Network (MCNN) to classify mango leaves infected by Anthracnose disease. They utilised a dataset of 1,070 images, comprising both healthy and infected leaves, captured at India’s Shri Mata Vaishno Devi University which is located in Katra, Jammu and Kashmir. The suggested MCNN model demonstrated higher classification accuracy of 97.13% compared to other state-of-the-art approaches.
@@
[Dwivedi2021] have put out the Grape Leaf Disease Detection Network (GLDDN) approach. Their proposed network uses Residual Convolution Neural Network (R-CNN) and multitask learning to identify esca, isariopsis, and black-rot in grape photos that are thoroughly explained in the proposed grape leaf disease detection network which is relies on multitasking, spatial and channel-wise attention processes, and Faster R-CNN for multi-level feature extraction. [Karthik2020] utilized a residual CNN enhanced with an attention mechanism to identify diseases in tomato leaves. An evaluation conducted on a dataset with 95,999 categorized tomato leaf images revealed a 98% accuracy rate for their approach. [Zeng2020] created a residual CNN with self-attention architecture for disease classification, achieving 98% accuracy on the MK-D2 dataset and an accuracy of 95.33% on the AES-CD9214 dataset. Researchers in [Chen2021SEMobileNet][CHEN2021] applied different attention strategies within well-known architectures such as MobileNet, Inception, and residual CNNs to identify crop diseases. While these approaches delivered strong results, their evaluation focused on only a limited selection of plant diseases, unlike broader models addressing multiple crops. [Cap2022] has proposed LeafGAN architecture for the augmentation of diseased leaf images via transformation for improving the plant disease diagnosis system using large scale dataset. Their architecture has own attention mechanism to transform only relevant areas from plant leaf images with different backgrounds, which ensures the diversity of training data. [garima_] proposed CROPCARE, which is a technology for detecting and preventing crop diseases in real time. This system integrates Computer Vision (CV), Google Cloud services, and the Internet of Things (IoT) for the detection of crop diseases in the plant village dataset. It uses the pre-trained model MobileNet-V2 and Super resolution Convolution Neural Network (SRCNN) for generating decisions of the model trained on various diseases. To solve the data imbalance problem of health vs unhealthy images [Zhao2022] have used a DoubleGAN architecture. Their architecture uses Wasserstein Generative Adversarial Networks (WGAN) and Superresolution Generative Adversarial Network (SRGAN) to balance the dataset. [PAL2023EAAI] has proposed a technique to categorize the severity of damaged plants, called Agriculture Detection (AgriDet). Their method combines traditional Inception-Visual Geometry Group Networks (INC-VGGN) with Kohonen-based deep learning networks. A solution for the smaller dataset is provided by [Liu2023] where they have proposed a network called PiTLiD. Their proposed network that uses a small sample size and pretrained Inception-V3 convolutional neural network and transfer learning to identify plant leaf diseases from plant leaf phenotypic data. Gehlot et al. [Gehlot2023MultimediaSys] has presented a new dataset for multiclass tomato specific disease which was evaluated using difference CNN architecture models. Zeng et al. [Zeng2023MultimediaSys] has presented a Squeeze-and-Excitation Vision Transformer (SEViT) which used ResNet with channel attention module for the classification of disease in different crop dataset crawled from different internet sources. Their results establishes the supremacy for many similar types of diseases.
The above-discussed methods rely on deep learning algorithms that operate as black boxes, making it challenging to comprehend how they make their decisions. While they achieve high accuracy, they lack interpretability, limiting trust and dependability in practical uses. These methods do not provide any explainable solution for deep learning models’ opaque nature that makes it difficult to understand their judgements. Therefore, explainability is necessary for trust, transparency, and validation, ensuring that model predictions are reliable and interpretable, especially in critical applications like plant disease detection. The introduction of XAI techniques, such as GradCAM and LRP, has addressed this issue by visualizing model decisions. [Arvind2021] developed a Computer-Aided Diagnostic System (CADS) incorporating an XAI framework to build trust in the system’s decisions. EfficientNet B5 outperformed VGG19, VGG16, Inception v3, ResNet, MobileNet, and other EfficientNet variants among the models used for transfer learning. The predictions made by EfficientNet-B5 were explained using XAI methods like LIME and Grad-CAM, with validation performed through YOLOv4. [Kinger2021] introduced a separate XAI-driven approach for using deep learning for disease detection. They found the illness and indicated the crucial leaf areas for the GradCAM++ explainer model’s classification. They began by utilising images from the Plant Village dataset of both healthy and damaged plant leaves. A pre-trained ImageNet system with a VGG16 deep learning model as a base was used by the authors. They have used GradCAM++’s visualisation technique for model interpretation, which effectively identified the afflicted areas and used heatmap overlays to highlight them. Although methods like GradCAM [Selvaraju2019], [Zhou2015] and LRP are commonly utilized to interpret CNN outputs, their application in detecting plant leaf diseases is still relatively new. GradCAM has shown success in areas like medical imaging, image categorisation, and object detection and has recently been explored for plant disease identification. Another widely used technique for analyzing deep neural networks is LRP, which assigns relevance scores backward from the output to the input features, providing detailed explanations of CNN decisions [binder2016layerwise]. In the context of plant disease identification, Layer-wise Relevance Propagation (LRP) has been utilized to highlight significant regions of a leaf, aiding in the detection of disease symptoms and affected zones [Tariq2024]. Nevertheless, the resulting visual outputs from LRP can be intricate and challenging to understand, which may hinder their usability in real-world applications. To improve interpretability, a number of studies have looked into combining GradCAM using additional methods of explanation, including Lime [Ribeiro2016], DeepLIFT [Shrikumar2019], and Shapley Value Sampling [Lundberg2017], with the goal of addressing GradCAM’s shortcomings, including its tendency to produce noisy or vague visual outputs. Despite progress, issues like unclear GradCAM explanations and the complexity of LRP visualizations remain unresolved. Integrating various explainable AI techniques offers a promising direction for improving model transparency and fostering trust and dependability in sensitive domains such as plant disease diagnosis. [MUKHERJEE2025EAAI] has presented an comprehensive review of rice disease detection techniques using AI based solutions. A detailed survey of recent advances in plant disease detection is provided by [Qadri2025]. They have highlighted the issue and challenges in plant leaf disease detection using machine learning and deep learning based solutions.
@@ 3 Proposed Methodology
The overview of interpretable architecture used for identifying plant leaf disease is provided in Fig. 1. The input plant images undergoes through data preprocessing to improve quality and guarantee alignment with the input structure of the proposed CBAM-VGG16 model. This involves histogram equalization to improve contrast, normalization of pixel values to the $[0,1]$ range and resizing images to $224\times 224$ pixels. Following preprocessing, the enhanced VGG16 [Simonyan2014] model integrated with CBAM is employed for disease classification. In this architecture, CBAM modules are added following every convolutional stage of the VGG16 model. The CBAM mechanism adaptively refines the feature maps by applying both channel and spatial attention, thereby enhancing the model’s ability to focus on disease-affected regions. Once the input image’s class label is predicted by the model, multiple explainability modules are used to interpret its decision-making process. CBAM attention maps, generated by the integrated CBAM modules in the VGG16 architecture, provide insights into the spatial and channel-wise focus of the network across layers. For class-discriminative localization, Grad-CAM [Selvaraju2019] and Grad-CAM++ [Chattopadhay2018] are used to generate class-discriminative heatmaps that draw attention to the input image’s most significant areas. Additionally LRP is employed using multiple propagation rules i.e. including $\epsilon$-rule, $\epsilon$-$\gamma$-box, $\alpha_{2}\beta_{1}$, and excitation backpropagation to assign pixel-level relevance scores, further enriching the interpretability of the model’s predictions through diverse attribution perspectives. This entire pipeline from data preprocessing to classification and visual explanation forms an end-to-end plant leaf disease detection system designed for improved interpretability and trust in AI-based agricultural diagnostics. The internal modules of the architecture are explained in further subsections.
3.1 Data Preprocessing
To improve the quality, consistency, and compatibility of the input images to the CBAM-enhanced VGG16 model, we apply a series of preprocessing steps, i.e., histogram equalization, normalisation, and resizing, visually presented in Fig 2.
3.1.1 Histogram Equalization
Histogram equalization using the Contrast Limited Adaptive Histogram Equalization (CLAHE) [pizer1990clahe] is used to improve the grape leaf picture’s contrast by redistributing pixel intensity values, as shown in Fig. 2. Unlike traditional histogram equalization, CLAHE operates on small image regions (tiles) and applies contrast enhancement while limiting noise amplification in homogeneous areas, making it more effective for preserving fine details in medical and agricultural images.
3.1.2 Normalization
Normalization scales pixel intensity values in the range $[0,1]$, ensuring uniformity across the dataset as illustrated in Fig. 2 This prevents biased learning caused by varying pixel ranges and improves model convergence and training stability.
3.1.3 Resizing
All images are resized to $224\times 224$ pixels, which matches the expected input size of the CBAM-VGG16 architecture. This resizing step maintains structural integrity while ensuring consistent dimensions for batch processing.
(a) Original Image
(b) Histogram Equalized
(c) Normalized Image
(d) Resized Image
Figure 2: Pre-processing of the input images
@@ 3.2 CBAM VGG16 Architecture
As illustrated in Fig. 1, the proposed CBAM-VGG16 architecture begins by utilising input layer of size $224\times 224\times 3$. In the architecture, CBAM layers are added after each MaxPooling2D layer. Even though CBAM is a lightweight architecture, overuse could lead to overfitting, which is why the CBAM layer is added after each MaxPooling layer. The channel module and the spatial module are the two primary parts of the CBAM developed by [woo2018cbamconvolutionalblockattention]. Given the input $I_{f}\in\mathbb{R}^{c\times W\times H}$, a 1D channel attention map is generated by CBAM $C_{m}\in\mathbb{R}^{c\times 1\times 1}$ as well as a 2D spatial attention map $S_{m}\in\mathbb{R}^{c\times W\times H}$, where $c$ in our scenario stands for three channel input, $W$ and $H$ stand for the input feature map’s width and height, respectively. Our suggested CBAM VGG16’s attention process is given in Eq. 1 and 2.
@@
where $\otimes$ denotes element-wise multiplication, $I_{c}f^{\prime}$ is the channel multiplied output, and $I_{s}f^{\prime\prime}$ denote the final refined output.
The placement of the CBAM layer after each MaxPooling layer is illustrated in Fig. 3. The internal workings of CBAM’s channel attention and spatial attention mechanisms are detailed in the following subsections.
Figure 3: Internal structure of CBAM layers
@@ 3.2.1 Channel Attention Module
The composition of the channel attention module is given in Fig. 4. A map of channel attention $C_{m}\in\mathbb{R}^{c\times 1\times 1}$ is applied to the input features by the channel attention module. To generate the average pooled feature, $I_{f}^{\text{avg}}$, and the max pooled feature, $I_{c}^{\text{max}}$, the average and max pooling procedures are applied independently. Thereafter, a shared Multilayer Perceptron (MLP) network with one hidden layer receives $I_{c}^{\text{avg}}$ and $I_{c}^{\text{max}}$. In our CBAM VGG16 design, the activation size of the hidden layer is set to $\mathbb{R}^{c/r\times 1\times 1}$, where the reduction ratio denoted by $r$, which is set to eight. This value provides offers an equitable trade-off between model complexity and representational capacity, as empirically validated in [woo2018cbamconvolutionalblockattention], ensuring sufficient channel inter dependencies are captured without incurring significant computational overhead. The channel attention method used in our proposed architecture is provided as given in Eq. 3 and 4.
$C_{m}(I_{f})$ $=\sigma\big(\text{MLP}(\text{AvgPool}(I_{f}))+\text{MLP}(\text{MaxPool}(I_{f}))\big)$ (3)
$C_{m}(I_{f})$ $=\sigma\big(W_{1}(W_{0}(I_{c}^{\text{avg}}))+W_{1}(W_{0}(I_{c}^{\text{max}}))\big)$ (4)
In the above equations, $\sigma$ denotes the sigmoid activation function. The weights $W_{0}\in\mathbb{R}^{c\times c/r}$ and $W_{1}\in\mathbb{R}^{c/r\times c}$ are shared parameters of the multi-layer perceptron (MLP). To generate the channel attention map $C_{m}(I_{f})$, the input feature map $I_{f}$ undergoes both average pooling and max pooling operations across its spatial dimensions, resulting in $I_{c}^{\text{avg}}$ and $I_{c}^{\text{max}}$, respectively. These are then passed through the shared MLP, and their outputs are summed and activated using the sigmoid function. The resulting channel attention map is element-wise multiplied with the original input $I_{f}$ to obtain the refined feature map $I_{c}f^{\prime}$, as defined in Eq. 1. The structure of the channel attention block in CBAM is depicted in Fig. 4.

### v2

@@ Abstract
Plant diseases pose a significant threat to global food security, necessitating accurate and interpretable disease detection methods. This study introduces an interpretable attention-guided Convolutional Neural Network (CNN), CBAM-VGG16, for plant leaf disease detection. By integrating Convolution Block Attention Module (CBAM) at each convolutional stage, the model enhances feature extraction and disease localization. Trained on five diverse plant disease datasets, our approach outperforms recent techniques, achieving high accuracy (up to 98.87%) and demonstrating robust generalization. Here, we show the effectiveness of our method through comprehensive evaluation and interpretability analysis using CBAM attention maps, Gradient-weighted Class Activation Mapping (Grad-CAM), Grad-CAM++, and Layer-wise Relevance Propagation (LRP). This study advances the application of explainable AI in agricultural diagnostics, offering a transparent and reliable system for smart farming.
@@
Plant leaf disease, Deep learning, Attention Enhanced CNN, Explainable AI, Agricultural Diagnostics, Smart Farming
@@ 1 Introduction
The agriculture industry is essential in maintaining food security worldwide, but different crops suffer from a variety of diseases due to varying weather conditions, posing a major threat to crop yield and quality. These diseases are often caused by factors such as extreme temperatures, microbial infections, and changes in humidity or soil conditions. Farmers have always relied on human examination to identify diseases, which is labour-intensive, prone to mistakes, and ineffective on large farms. As a result, among the most crucial areas of research in smart agriculture is the automation of plant disease identification and classification utilizing various Artificial Intelligence (AI) approaches [Singh2019, Dwivedi2021]. In recent years, Convolutional Neural Networks (CNNs) based models in agriculture have performed well on challenges such as classifying plant diseases. However, despite their effectiveness, CNNs often operate as black boxes, offering limited transparency into how predictions are made, hampering trust and broader adoption.
@@
In order to tackle this, Explainable AI (XAI) methods have been developed enabling visual interpretation of model decisions through attention maps, GradCAM [Selvaraju2017], and LRP [Bach2015]. While GradCAM and GradCAM++ [Chattopadhay2018] generate class-discriminative localization maps using gradients from the final convolutional layers, they may lack resolution and be susceptible to noisy activations. LRP, in contrast, provides pixel-level attributions by backpropagating the model output through a set of layer-specific relevance propagation rules. This offers a more granular explanation of predictions, which is crucial in medical and agricultural diagnostics.
@@
In this work, an explainable deep learning approach is proposed for the detection of plant leaf disease. Our architecture is based on the VGG16 [Simonyan2014] backbone enhanced with the CBAM [woo2018cbamconvolutionalblockattention] which introduces attention layers for inherent interpretability of the model with an emphasis on the most relevant features at both the channel and spatial levels. Following each of the five convolutional layers, CBAM modules are added to improve classification accuracy and localization of relevant features by capturing both spatial and channel-wise attention. Five distinct datasets are used to train the model, namely Apple, Plant Village, Embrapa, Maize and Rice to ensure the generalizability and applicability of our proposed method across diverse set of crops. Apart from the inherent interpretability of the proposed method’s decision-making process through CBAM layers, we also demonstrate interpretability using advanced explainability methods like LRP, Grad-CAM, and Grad-CAM++. We have also employed high-dimensional feature visualization techniques such as t-distributed Stochastic Neighbor Embedding (t-SNE) and Uniform Manifold Approximation and Projection (UMAP) to visualize feature in lower dimension for visualization of extracted features. Overall, this study advances the application of XAI for agricultural use by presenting an interpretable and performance-robust framework for plant disease classification. The following are thes main contributions of the proposed work.
@@
  • Superior performance as compared with other state-of-the-art methods on all five datasets demonstrates the generalization and adaptive capability of proposed method for any crop leaf disease identification.
@@ 2 Related Works
Barbedo [Barbedo2018] applied CNN architectures to analyze how variations in samples and dataset size influence the efficiency of identifying plant diseases via transfer learning. For this purpose, a dataset comprising 1,383 background-free images across 56 distinct disease types was compiled. Singh et al. [Singh2019] developed a Multilayer Convolutional Neural Network (MCNN) to classify mango leaves infected by Anthracnose disease. The suggested MCNN model demonstrated higher classification accuracy of 97.13% compared to other state-of-the-art approaches. Dwivedi et al. [Dwivedi2021] have put out the Grape Leaf Disease Detection Network (GLDDN) approach. Their proposed network uses Residual Convolution Neural Network (R-CNN) and multitask learning to identify esca, isariopsis, and black-rot in grape photos. Researchers in [CHEN2021] [Chen2021SEMobileNet] applied different attention strategies within well-known architectures such as MobileNet, Inception, and residual CNNs to identify crop diseases. While these approaches delivered strong results, their evaluation focused on only a limited selection of plant diseases, unlike broader models addressing multiple crops. Cap et al. [Cap2022] has proposed LeafGAN architecture for the augmentation of diseased leaf images via transformation for improving the plant disease diagnosis system using large scale dataset. To solve the data imbalance problem of health vs unhealthy images Zhao et al.[Zhao2022] have used a DoubleGAN architecture. Their architecture uses Wasserstein Generative Adversarial Networks (WGAN) and Superresolution Generative Adversarial Network (SRGAN) to balance the dataset. Some of the recent methods also used the fuzzy rank-based ensemble along with the pre-trained CNNs [Medjadba2025] and fuzzy feature extraction [Nagi2023] for the plant leaf detection.
@@
Although methods like GradCAM [Selvaraju2019] and LRP are commonly utilised to interpret CNN outputs, their application in detecting plant leaf diseases is still relatively new. In the context of plant disease identification, LRP has been utilised to highlight significant regions of a leaf, aiding in the detection of disease symptoms and affected zones [Tariq2024]. To improve interpretability, a number of studies have looked into combining GradCAM using additional methods of explanation with the goal of addressing GradCAM’s shortcomings, including its tendency to produce noisy or vague visual outputs. Despite progress, issues like unclear GradCAM explanations and the complexity of LRP visualisations remain unresolved. A detailed survey of recent advances in plant disease detection is provided by Qadri et al. [Qadri2025]. They have highlighted the issue and challenges in plant leaf disease detection using machine learning and deep learning based solutions.
@@ 3 Proposed Methodology
The overview of interpretable architecture used for identifying plant leaf disease is provided in Fig. 1. The input plant images undergo through data preprocessing to improve quality and guarantee alignment with the input structure of the proposed CBAM-VGG16 model. This involves Contrast Limited Adaptive Histogram Equalization (CLAHE) to improve contrast, normalization of pixel values to the $[0,1]$ range and resizing images to $224\times 224$ pixels. Following preprocessing, the enhanced VGG16 [Simonyan2014] model integrated with CBAM is employed for disease classification. In this architecture, CBAM modules are added following every convolutional stage of the VGG16 model. The CBAM mechanism adaptively refines the feature maps by applying both channel and spatial attention, thereby enhancing the model’s ability to focus on disease-affected regions. Once the input image’s class label is predicted by the model, multiple explainability modules are used to interpret its decision-making process. CBAM attention maps, generated by the integrated CBAM modules in the VGG16 architecture, provide insights into the spatial and channel-wise focus of the network across layers. For class-discriminative localization, Grad-CAM [Selvaraju2019] and Grad-CAM++ [Chattopadhay2018] are used to generate class-discriminative heatmaps that draw attention to the input image’s most significant areas. Additionally LRP is employed using multiple propagation rules i.e. including $\epsilon$-rule, $\epsilon$-$\gamma$-box, $\alpha_{2}\beta_{1}$, and excitation backpropagation to assign pixel-level relevance scores, further enriching the interpretability of the model’s predictions through diverse attribution perspectives. This entire pipeline from data preprocessing to classification and visual explanation forms an end-to-end plant leaf disease detection system designed for improved interpretability and trust in AI-based agricultural diagnostics. The internal modules of the architecture are explained in further subsections.
@@ 3.1 CBAM VGG16 Architecture
As illustrated in Fig. 1, the proposed CBAM-VGG16 architecture begins by utilising input layer of size $224\times 224\times 3$. In the architecture, CBAM layers are added after each MaxPooling2D layer. Even though CBAM is a lightweight architecture, overuse could lead to overfitting, which is why the CBAM layer is added after each MaxPooling layer. The channel module and the spatial module are the two primary parts of the CBAM developed by [woo2018cbamconvolutionalblockattention]. Given the input $I_{f}\in\mathbb{R}^{c\times W\times H}$, a 1D channel attention map is generated by CBAM $C_{m}\in\mathbb{R}^{c\times 1\times 1}$ as well as a 2D spatial attention map $S_{m}\in\mathbb{R}^{c\times W\times H}$, where $c$ in our scenario stands for three channel input, $W$ and $H$ stand for the input feature map’s width and height, respectively. Our suggested CBAM VGG16’s attention process is given in Eqs. 1 and 2.
@@
where $\otimes$ denotes element-wise multiplication, $I_{c}f^{\prime}$ is the channel multiplied output, and $I_{s}f^{\prime\prime}$ denotes the final refined output. The internal workings of CBAM’s channel attention and spatial attention mechanisms are detailed in the following subsections.
@@ 3.1.1 Channel Attention Module
A map of channel attention $C_{m}\in\mathbb{R}^{c\times 1\times 1}$ is applied to the input features by the channel attention module. To generate the average pooled feature, $I_{f}^{\text{avg}}$, and the max pooled feature, $I_{c}^{\text{max}}$, the average and max pooling procedures are applied independently. Thereafter, a shared Multi Layer Perceptron (MLP) network with one hidden layer receives $I_{c}^{\text{avg}}$ and $I_{c}^{\text{max}}$. In our CBAM VGG16 design, the activation size of the hidden layer is set to $\mathbb{R}^{c/r\times 1\times 1}$, where the reduction ratio denoted by $r$, which is set to eight. This value provides an equitable trade-off between model complexity and representational capacity, as empirically validated in [woo2018cbamconvolutionalblockattention], ensuring sufficient channel interdependencies are captured without incurring significant computational overhead. The channel attention method used in our proposed architecture is provided as given in Eqs. 3 and 4.
$\footnotesize C_{m}(I_{f})$ $=\sigma\big(\text{MLP}(\text{AvgPool}(I_{f}))+\text{MLP}(\text{MaxPool}(I_{f}))\big)$ (3)
$\footnotesize C_{m}(I_{f})$ $=\sigma\big(W_{1}(W_{0}(I_{c}^{\text{avg}}))+W_{1}(W_{0}(I_{c}^{\text{max}}))\big)$ (4)
In the above equations, $\sigma$ denotes the sigmoid activation function. The weights $W_{0}\in\mathbb{R}^{c\times c/r}$ and $W_{1}\in\mathbb{R}^{c/r\times c}$ are shared parameters of the MLP. To generate the channel attention map $C_{m}(I_{f})$, the input feature map $I_{f}$ undergoes both average pooling and max pooling operations across its spatial dimensions, resulting in $I_{c}^{\text{avg}}$ and $I_{c}^{\text{max}}$, respectively. These are then passed through the shared MLP, and their outputs are summed and activated using the sigmoid function.